# The 2-systole on compact Kähler surfaces with positive scalar curvature

### v1

@@ Abstract.
We study the 2-systole on compact Kähler surfaces of positive scalar curvature. For any such surface $(X,\omega)$, we prove the sharp estimate $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 12\pi$, with equality if and only if $X=\mathbb{P}^{2}$ and $\omega$ is the Fubini–Study metric. Using the classification of positive scalar curvature Kähler surfaces by their minimal models, we also determine the optimal constant in each case and describe the corresponding rigid models: $12\pi$ when the minimal model is $\mathbb{P}^{2}$, $8\pi$ for Hirzebruch surfaces, and $4\pi$ for non-rational ruled surfaces. In the non-rational ruled case, we also give an independent analytic proof, adapting Stern’s level set method to the holomorphic fibration in Kähler setting.
@@ 1. Introduction
Systolic geometry studies how topology constrains the existence of low-volume representatives of nontrivial homology classes, and how these minimal “sizes” interact with global geometric constraints. Following Berger’s terminology [Ber72a, Ber72b, Ber81] (see also [Ber08]) and Gromov’s subsequent development of systolic geometry [Gro83], if $(M,g)$ is a closed Riemannian manifold of dimension $n\geq k$, the kk–systole of $(M,g)$ is defined by
@@
In the setting of positive scalar curvature (PSC), a particularly natural object is the 2-systole. Indeed, scalar curvature is obtained by averaging sectional curvatures over $2$-planes, and a basic mechanism for producing PSC metrics is to insert an $S^{2}$-component: for product metrics, one has $\text{scal}(g\oplus h)=\text{scal}(g)+\text{scal}(h)$, so $S^{2}(r)\times(N,h)$ carries PSC metrics once $r$ is sufficiently small, to a large extent, independently of the geometry of $N$. This illustrates that scalar curvature positivity is often controlled by the $2$-sphere direction, whose size is measured by the $2$-systole.
In dimension three, the interplay between $2$-systolic quantities and PSC is by now well understood. If $(M^{3},g)$ is PSC, Schoen-Yau [SY79] proved that any area–minimizing surface in $M$ is homeomorphic to either $S^{2}$ or $\mathbb{RP}^{2}$, showing that the relevant minimizing surfaces are essentially spherical. Building on this, Bray-Brendle-Neves [BBN10] established a sharp $\pi_{2}$-systolic inequality. Denote by $\operatorname{sys}^{\pi_{2}}_{2}(g)$ the infimum of the areas of homotopically nontrivial 22–spheres in $(M^{3},g)$. Then
@@
with equality if and only if the universal cover of $(M^{3},g)$ is isometric to the Riemannian product $S^{2}\times\mathbb{R}$ endowed with the round metric on $S^{2}$ and the flat metric on $\mathbb{R}$. In [Ste22], Stern gave a new proof of (1.1) leveraging the level set method. The bound (1.1) was recently refined by a quantitative gap theorem of Xu [Xu25], and there has been substantial further progress on systolic inequalities under scalar curvature assumptions; see for instance [Zhu20, Ric20, Ori25].
@@
$\operatorname{sys}_{2}(\omega):=\operatorname{sys}_{2}(g_{\omega}),$
while $S(\omega)$ denotes the (Chern) scalar curvature of $\omega$ (in the Kähler setting, we have $2S(\omega)=\text{scal}(g_{\omega})$). With this convention, all our main inequalities are stated purely in terms of the Kähler form $\omega$. Our result shows that:
@@
It is worth noting that a compact Kähler surface $X$ admits a PSC metric if and only if $X$ is obtained from $\mathbb{P}^{2}$ or a ruled surface $\mathbb{P}(\mathcal{E})$ by a finite sequence of blow-ups. More precisely, Yau showed that the existence of a Kähler metric with positive total scalar curvature forces the Kodaira dimension of $X$ to be -$\infty$ [Yau74], and the Enriques-Kodaira classification then implies that any minimal compact Kähler surface with PSC is either $\mathbb{P}^{2}$ or ruled [BHPVdV04, Chapter V]. Building on this and using Seiberg-Witten theory, LeBrun proved that for minimal complex surfaces of Kähler type, the existence of a Riemannian metric of positive scalar curvature is equivalent to the existence of a Kähler PSC metric, and that this happens if and only if the surface is either $\mathbb{P}^{2}$ or ruled [LeB95, Theorem 4]. He conjectured that the same characterization remains true after allowing blow-ups. The remaining gap, whether blowing up preserves the sign of the scalar curvature, was settled more recently by Brown, who proved that blow-ups with small data preserve the sign of scalar curvature and thereby completed the classification [Bro24, Theorem B]. In particular, every PSC Kähler surface has minimal model either $\mathbb{P}^{2}$ or a ruled surface. It is therefore natural to restate Theorem 1.1 in a form that records how the optimal constant depends on the minimal model. This leads to the following three-way refinement.
@@
    (Theorem 3.4) If the minimal model of $X$ is $\mathbb{P}^{2}$, then
@@
    (Theorem 4.10) If the minimal model of $X$ is a Hirzebruch surface $\mathbb{F}_{e}$ fibred over $\mathbb{P}^{1}$, then
@@
    (Theorems 4.5 and 5.4) If the minimal model of $X$ is a ruled surface fibred over a base curve $B$ of genus $g(B)\geq 1$, then
@@
    with equality if and only if $B$ is an elliptic curve, the universal cover of $X$ is biholomorphic to $\mathbb{P}^{1}\times\mathbb{C}$, and $\omega$ is induced by the product of the Fubini–Study metric on $\mathbb{P}^{1}$ with a flat metric so that $\operatorname{sys}_{2}(\omega)$ is realized by a $\mathbb{P}^{1}$–fibre.
@@
A key feature of the Kähler setting is that several genuinely Riemannian quantities can be expressed directly in terms of the Kähler class. In real dimension four, complex curves are calibrated by the Kähler form: if $(X,\omega)$ is a compact Kähler manifold and $C\subset X$ is a smooth complex curve, then
$\operatorname{Area}_{\omega}(C)=\int_{C}\omega=[\omega]\cdot[C],$
so the $\omega$–area of $C$ is entirely encoded by the cohomology class $\omega]\in H^{1,1}(X)$ and the homology class $C]\in H_{2}(X;\mathbb{Z})$. More generally, if $C$ is an effective (possibly singular or reducible) curve, we view $C=\sum_{i}m_{i}C_{i}$ as a positive integral real 22–cycle, and the pairing can be understood as
@@
Motivated by this interplay between geometry and intersection theory, we introduce the following algebraic analogue of the 22–systole.
@@
Let $(X,\omega)$ be a compact Kähler manifold. The holomorphic $2$-systole of a Kähler class $\omega]$ is defined by
@@
In other words, $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$ measures the smallest possible $\omega$-area of a nontrivial effective curve on $X$, expressed purely in terms of the intersection pairing. We emphasize that there exist compact Kähler manifolds carrying no complex curves (for instance, very general K3 surfaces with Picard number 0 or non-algebraic complex tori), in which case the set of candidates in (1.2) is empty. Hence, Definition 1.3 is worth using only when $X$ carries at least one effective curve. Fortunately, in the presence of positive scalar curvature, this is automatic (and $\operatorname{sys}^{\operatorname{hol}}_{2}$ is finite): by [Yan19, Theorem 1.3], a compact complex manifold admits a Hermitian metric with positive Chern scalar curvature if and only if $K_{X}$ is not pseudo-effective. If moreover $X$ is Kähler, then by Ou’s characterization [Ou25, Theorem 1.1] this is equivalent to $X$ being uniruled (i.e. covered by rational curves). Summarizing,
[…]
and hence $X$ carries plenty of effective curves, so the infimum in (1.2) is taken over a nonempty set. This parallels the geometric intuition from the Riemannian setting: scalar curvature positivity is often driven by the presence of 2-sphere components $S^{2}\cong\mathbb{P}^{1}$.
We then consider the following scale-invariant functional on the Kähler cone $\mathcal{K}(X)$:
$\mathcal{J}_{X}([\omega]):=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\hat{S}([\omega])=\left(\inf_{C\ \mathrm{eff.}}[\omega]\cdot[C]\right)\cdot\frac{2n\pi\,c_{1}(X)\cup[\omega]^{n-1}}{[\omega]^{n}},$
where $\hat{S}([\omega])$ is the (normalized) average Chern scalar curvature, $n=\dim_{\mathbb{C}}X$ and $c_{1}(X)$ is the first Chern class. Since holomorphic curves are admissible competitors for the homological 22–systole and are calibrated by $\omega$, one always has
$\operatorname{sys}_{2}(\omega)\;\leq\;\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]).$
Moreover, $\min_{X}S(\omega)\leq\hat{S}([\omega])$, hence
@@
Equality can occur only if $S(\omega)$ is constant (i.e. $\omega$ is cscK) and the $2$-systole is realized by a holomorphic curve (as a calibrated cycle).
From this point on, we restrict to Kähler surfaces. In this case
$\mathcal{J}_{X}([\omega])=\left(\inf_{C\ \mathrm{eff.}}[\omega]\cdot[C]\right)\cdot\frac{4\pi\,c_{1}(X)\cdot[\omega]}{[\omega]^{2}}$
is completely determined by intersection numbers of the Kähler class $\omega]$ with curve classes in $H_{2}(X;\mathbb{Z})$ and with $c_{1}(X)$. In particular, every PSC Kähler surface is projective (indeed, in complex dimension $2$ the PSC classification forces the minimal model to be rational or ruled, hence projective, and projectivity is preserved under blow-ups). On a projective surface, the Nakai–Moishezon criterion implies that a real (1,1)(1,1)–class is Kähler if and only if
@@
$\omega]^{2}>0\quad\text{and}\quad[\omega]\cdot C>0\ \text{for every irreducible curve }C\subset X.$
Since we are interested in PSC Kähler metrics, we further restrict to classes with $c_{1}(X)\cdot[\omega]>0$. The global inequality in Theorem 1.1 then reduces to obtaining sharp upper bounds on $\sup\mathcal{J}_{X}$ over the slice
$\{[\omega]\in\mathcal{K}(X)\mid c_{1}(X)\cdot[\omega]>0\},$
and, as we show in Theorem 2.2, this supremum is always finite. In particular, $\mathcal{J}_{X}$ does not blow up along sequences of Kähler classes approaching the boundary of the Kähler cone.
The remaining task is therefore a purely algebraic optimization problem on the Kähler cone of blow-ups. Let $\pi\colon X_{k}\to X$ be the blow-up of a minimal surface $X$ at $k$ points in very general position. Any Kähler class on $X_{k}$ can be written as
$\omega_{t}]=\pi^{*}[\omega]-\sum_{i=1}^{k}t_{i}E_{i}$
for some $\omega]\in\mathcal{K}(X)$ and some small coefficients $t_{i}>0$. Since $c_{1}(X_{k})=\pi^{*}c_{1}(X)-\sum_{i=1}^{k}E_{i}$, after fixing $\omega]$ we may view $\mathcal{J}_{X_{k}}([\omega_{t}])$ as a multivariable function of $t=(t_{1},\dots,t_{k})$. The main difficulty is to understand the geometric quantity
$m(t):=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}]).$
When the Mori cone $\overline{\mathrm{NE}}(X_{k})$ is rational polyhedral and generated by finitely many effective curves (for instance, for del Pezzo surfaces), one can, in principle, enumerate its extremal rays and reduce the problem to a finite-dimensional optimization (see Examples 3.5 and 4.9). The main difficulty arises for blow-ups with infinitely generated Mori cone, where such an enumeration is impossible.
To overcome this, we start from the identity
$\mathcal{J}_{X_{k}}([\omega_{t}])=4\pi\,m(t)\cdot\frac{c_{1}(X)\cdot[\omega]-\sum_{i=1}^{k}t_{i}}{[\omega]^{2}-\sum_{i=1}^{k}t_{i}^{2}}.$
Set $S(t):=\sum_{i=1}^{k}t_{i}$ and $Q(t):=\sum_{i=1}^{k}t_{i}^{2}$. Since each $E_{i}$ is an effective curve, we always have the coarse lower bounds $t_{i}\geq m(t)$ for all $i$, hence $S(t)\geq k\,m(t)$. Fix $(m,S)$, we consider the purely numerical family of vectors $t$ satisfying only
$t_{i}\geq m,\qquad\sum_{i=1}^{k}t_{i}=S.$
On this set, the factor $m\cdot\bigl(c_{1}(X)\cdot[\omega]-S\bigr)$ is fixed and the function $Q\mapsto 1/([\omega]^{2}-Q)$ is increasing. Therefore, for any geometric class $\omega_{t}]$ with numerical data $(m(t),S(t))$ we obtain the upper bound
$\mathcal{J}_{X_{k}}([\omega_{t}])\leq 4\pi\,\phi_{k}\bigl(m(t),S(t)\bigr),$
where
$\phi_{k}(m,S):=m\cdot\frac{c_{1}(X)\cdot[\omega]-S}{[\omega]^{2}-Q_{\max}(m,S)},\qquad Q_{\max}(m,S):=\max\left\{\sum_{i=1}^{k}t_{i}^{2}\ \Big|\ t_{i}\geq m,\ \sum t_{i}=S\right\}.$
By Proposition 3.1, the maximum $Q_{\max}(m,S)$ is achieved at the extremal vector
$t^{*}=(m,\dots,m,\ S-(k-1)m),$
so that
$Q_{\max}(m,S)=(k-1)m^{2}+\bigl(S-(k-1)m\bigr)^{2}.$
In particular, the bound $\mathcal{J}_{X_{k}}([\omega_{t}])\leq 4\pi\,\phi_{k}\bigl(m(t),S(t)\bigr)$ depends on $t$ only through $(m(t),S(t))$.
It remains to control the possible values of $(m(t),S(t))$. Instead of using all effective curve classes (which is infeasible when $\overline{\mathrm{NE}}(X_{k})$ is infinitely generated), we choose a finite collection of effective curves $C$ on $X_{k}$ and use the elementary inequalities $m(t)\leq[\omega_{t}]\!\cdot C$ to obtain explicit constraints on $(m(t),S(t))$. These finitely many constraints define a coarse admissible region $\mathcal{D}_{k}$ for $(m,S)$, which contains the true set of geometric data arising from Kähler classes. Consequently,
$\sup_{\mathcal{K}(X_{k})}\mathcal{J}_{X_{k}}([\omega_{t}])\leq 4\pi\,\sup_{(m,S)\in\mathcal{D}_{k}}\phi_{k}(m,S),$
turning an a priori infinite collection of curve constraints into a manageable finite-dimensional optimization problem. The details are worked out first for blow-ups of $\mathbb{P}^{2}$, and then adapted to Hirzebruch surfaces.
Finally, for non-rational PSC Kähler surfaces (i.e. ruled surfaces over bases of genus $g\geq 1$) we show that the same strategy applies. In addition, we develop an independent analytic approach inspired by Stern’s level-set method for harmonic maps [Ste22], adapted to the holomorphic fibration in the Kähler setting. Both the algebraic and analytic approaches lead to the same optimal constant and the same rigid model in Theorem 1.2 (3).
@@
It is instructive to compare the preceding discussion with the purely Riemannian PSC setting in higher dimensions. On a closed Riemannian manifold $(M^{n},g)$, a homological $k$-systole $\operatorname{sys}_{k}(g)$ is only defined (and finite) when $H_{k}(M;\mathbb{Z})\neq 0$. Otherwise the notion is vacuous. Even when $H_{k}(M;\mathbb{Z})\neq 0$, one should not expect any meaningful upper control of $\operatorname{sys}_{k}(g)$ from a pointwise lower bound $\text{scal}(g)\geq\sigma>0$ in general. Indeed, for $n\geq 3$ and $1\leq k\leq n-2$ with $k\neq n/2$ when $n$ is even, consider
$M=S^{k}\times S^{n-k},\qquad g_{t}:=t^{2}g_{S^{k}}\oplus g_{S^{n-k}},\qquad t\geq 1,$
where $g_{S^{m}}$ is the round metric of sectional curvature $1$. Then $\text{scal}(g_{t})\geq(n-k)(n-k-1)>0$, while $H_{k}(M;\mathbb{Z})\cong\mathbb{Z}$ is generated by $S^{k}\times\{\ast\}]$, and a standard argument shows that any nontrivial $k$-cycle has volume $\mathcal{O}(t^{k})$. Hence $\operatorname{sys}_{k}(g_{t})\to\infty$ as $t\to\infty$. [^1] The middle-dimensional case may behave differently even for simple products, e.g. on $S^{m}\times S^{m}$ one has $H_{m}\cong\mathbb{Z}^{2}$ and the naive one-sided scaling does not force $\operatorname{sys}_{m}$ to diverge.
This pathology has no Kähler analogue in the basic model. For instance, on $\mathbb{P}^{n}$, for every $1\leq k\leq n-1$ one has $H_{2k}(\mathbb{P}^{n};\mathbb{Z})\cong\mathbb{Z}$, generated by the class of a linear subspace $\mathbb{P}^{k}\subset\mathbb{P}^{n}$. Thus, unlike the sphere $S^{n}$ (whose homology vanishes in all intermediate degrees), the natural even-dimensional systolic questions on $\mathbb{P}^{n}$ are a priori non-vacuous, and the simplest PSC counterexamples based on “inflating one factor while keeping another PSC factor fixed” do not apply to Kähler manifolds such as $\mathbb{P}^{n}$.
More importantly, on a compact PSC Kähler manifold, the presence of 22–sphere components is not accidental but structural. As we illustrated before, PSC Kähler manifolds must be uniruled, which carry “plenty” of $\mathbb{P}^{1}$. Therefore, the $2$-systole (and, more generally, $2k$-systole via effective cycles when available) becomes a natural invariant on PSC Kähler manifold. This is the conceptual reason why high-dimensional systolic inequalities are more plausible in the PSC Kähler world, and it motivates our focus on the holomorphic $2$-systole and the scale-invariant functional $\mathcal{J}_{X}$.
@@
In Section 2 we review the basic notions used throughout the paper. We also show that, for projective Kähler surfaces in our setting, the functional $\mathcal{J}_{X}([\omega])$ is well-defined and satisfies $\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}<\infty$.
@@
In Section 4 we study ruled surfaces from an algebro-geometric perspective. We first treat the non-rational ruled case, which is technically simpler in our approach, and show that $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 4\pi$, together with the corresponding rigid model. We then recall the geometry of Hirzebruch surfaces and their blow-ups, and adapt the mass-shift argument to the rational ruled case. This leads to the sharp bound $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 8\pi$ for all PSC Kähler surfaces with Hirzebruch surfaces as minimal model, with rigidity characterized by the product Fubini-Study metric on $\mathbb{P}^{1}\times\mathbb{P}^{1}$.

### v2

@@ Abstract.
We study the 2-systole on compact Kähler surfaces of positive scalar curvature. For any such surface $(X,\omega)$, we prove the sharp estimate $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 12\pi$, with equality if and only if $X=\mathbb{P}^{2}$ and $\omega$ is the Fubini–Study metric. Using the classification of positive scalar curvature Kähler surfaces, we determine the optimal constant in each case and describe the corresponding rigid models. When $X$ is a non-rational ruled surface, we also give an independent analytic proof, adapting Stern’s level set method to the holomorphic fibration in Kähler setting.
@@ 1. Introduction
Systolic geometry studies quantitative lower bounds for nontrivial homology classes represented by cycles of minimal volume. Following Berger’s terminology [4, 3, 5] (see also [6]) and Gromov’s subsequent development of systolic geometry [12], if $(M,g)$ is a closed Riemannian manifold of dimension $n\geq k$, the kk–systole of $(M,g)$ is defined by
@@
In the setting of positive scalar curvature (PSC), a particularly natural object is the 2-systole. Indeed, scalar curvature is obtained by averaging sectional curvatures over $2$-planes, and a basic mechanism for producing PSC metrics is to insert an $S^{2}$-component: for product metrics, one has $\text{scal}(g\oplus h)=\text{scal}(g)+\text{scal}(h)$, so $S^{2}(r)\times(N,h)$ carries PSC metrics once $r$ is sufficiently small, to a large extent, independently of the geometry of $N$. This suggests that, in many PSC constructions, the relevant geometric scale is encoded by a $2$-sphere direction, and hence by a $2$-systolic quantity.
In dimension three, the interplay between $2$-systole and PSC is by now well understood. If $(M^{3},g)$ is PSC, Schoen-Yau [21] proved that any area–minimizing surface in $M$ is homeomorphic to either $S^{2}$ or $\mathbb{RP}^{2}$, showing that area-minimizing surfaces in the PSC setting are forced to have spherical topology. Building on this, Bray-Brendle-Neves [7] established a sharp $\pi_{2}$-systolic inequality. Denote by $\operatorname{sys}^{\pi_{2}}_{2}(g)$ the infimum of the areas of homotopically nontrivial 22–spheres in $(M^{3},g)$. Then
@@
with equality if and only if the universal cover of $(M^{3},g)$ is isometric to the Riemannian product $S^{2}\times\mathbb{R}$ endowed with the round metric on $S^{2}$ and the flat metric on $\mathbb{R}$. In [22], Stern gave a new proof of (1.1) leveraging the level set method. The bound (1.1) was recently refined by a quantitative gap theorem of Xu [23], and there has been substantial further progress on systolic inequalities under scalar curvature assumptions; see for instance [26, 20, 18].
@@
$\operatorname{sys}_{2}(\omega):=\operatorname{sys}_{2}(g_{\omega}).$
We also write $S(\omega)$ for the Chern scalar curvature of $\omega$; for Kähler metrics, one has $2S(\omega)=\text{scal}(g_{\omega})$), so positivity of $S(\omega)$ is equivalent to positive Riemannian scalar curvature. With this convention, all our main inequalities are stated purely in terms of the Kähler form $\omega$.
@@
The birational classification of PSC Kähler surfaces is by now complete. Yau proved that a compact Kähler surface with positive total scalar curvature has Kodaira dimension -$\infty$ [25]. By the Enriques–Kodaira classification, a minimal such surface is therefore either $\mathbb{P}^{2}$ or ruled [2, Chapter V]. In addition, LeBrun showed that for minimal complex surfaces of Kähler type, the existence of a PSC Riemannian metric is equivalent to the existence of a PSC Kähler metric, and this occurs precisely for $\mathbb{P}^{2}$ and ruled surfaces [17, Theorem 4]. The remaining birational step was completed by Brown, who proved that blow-ups preserve the sign of scalar curvature for sufficiently small data [8, Theorem B]. Thus a compact Kähler surface admits a PSC Kähler metric if and only if it is birational to $\mathbb{P}^{2}$ or to a ruled surface. It is therefore natural to restate Theorem 1.1 in a form that records how the optimal constant depends on the classification. This leads to the following three-way refinement.
@@
    (Theorem 3.4) If $X$ is birational to $\mathbb{P}^{2}$, then
@@
    (Theorem 4.11) If $X$ is a rational ruled surface fibred over $\mathbb{P}^{1}$, then
@@
    (Theorems 4.5 and 5.4) If $X$ is a non-rational ruled surface fibred over a base curve $B$ of genus $g(B)\geq 1$, then
@@
    with equality if and only if $B$ is an elliptic curve, the universal cover of $X$ is biholomorphic to $\mathbb{P}^{1}\times\mathbb{C}$, and $\omega$ is induced by the product of the Fubini–Study metric on $\mathbb{P}^{1}$ and a flat metric on $\mathbb{C}$, with $\operatorname{sys}_{2}(\omega)$ is realized by a $\mathbb{P}^{1}$–fibre.
@@
A key feature of the Kähler setting is that several quantities relevant to our problem admit a direct cohomological interpretation. In real dimension four, complex curves are calibrated by the Kähler form: if $(X,\omega)$ is a compact Kähler surface and $C\subset X$ is a smooth complex curve, then
$\operatorname{Area}_{\omega}(C)=\int_{C}\omega=[\omega]\cdot[C].$
More generally, if $C=\sum_{i}m_{i}C_{i}$ is an effective (possibly singular or reducible) curve, viewed as a positive integral real 22–cycle, then
@@
Thus the $\omega$–area of an effective curve is encoded entirely by the Kähler class $\omega]$ and the corresponding curve class in homology. This motivates the following algebraic analogue of the 22–systole.
@@
Let $(X,\omega)$ be a compact Kähler manifold. The holomorphic 22–systole of a Kähler class $\omega]$ is defined by
@@
In general, the infimum in (1.2) need not be taken over a nonempty set, since a compact Kähler manifold may contain no effective curves. In the PSC Kähler setting, however, this issue does not arise. By [24, Theorem 1.3], a compact complex manifold admits a Hermitian metric with positive Chern scalar curvature if and only if its canonical bundle is not pseudo-effective. If moreover the manifold is Kähler, then by [19, Theorem 1.1] this is equivalent to uniruledness. In particular, every compact PSC Kähler manifold contains plenty of effective curves, so $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])$ is well-defined and finite in the setting of this paper.
We then introduce the scale-invariant functional
$\mathcal{J}_{X}([\omega]):=\operatorname{sys}^{\operatorname{hol}}_{2}([\omega])\cdot\hat{S}([\omega]),$
where $\hat{S}([\omega])$ denotes the normalized average Chern scalar curvature of the Kähler class $\omega]$. Since this paper concerns Kähler surfaces, we may write
The importance of $\mathcal{J}_{X}$ comes from the elementary comparison
$\operatorname{sys}_{2}(\omega)\leq\operatorname{sys}^{\operatorname{hol}}_{2}([\omega]),$
because every effective curve gives a nontrivial real 22–cycle, together with the inequality
$\min_{X}S(\omega)\leq\hat{S}([\omega]).$
Hence
@@
Therefore, the PSC $2$-systolic inequalities in this paper follow from sharp upper bounds for $\mathcal{J}_{X}$. Moreover, equality in the above comparison can occur only if $\omega$ has constant scalar curvature and $\operatorname{sys}_{2}(\omega)$ is realized by a holomorphic curve.
For PSC Kähler surfaces, this becomes a problem in intersection theory. Indeed, such surfaces are projective, and the Kähler cone admits a numerical description via the Nakai–Moishezon criterion. We may therefore regard $\mathcal{J}_{X}$ as a function on the cone of classes $\omega]$ satisfying
@@
$\omega]^{2}>0,\qquad[\omega]\cdot C>0\ \text{for every irreducible curve }C\subset X,\qquad c_{1}(X)\cdot[\omega]>0.$
A first step, carried out in Section 2, is to show that $\mathcal{J}_{X}$ remains uniformly bounded on this region.
The main difficulty appears on blow-ups. If $\pi\colon X_{k}\to X$ is the blow-up of a minimal surface $X$ at $k$ points, then any Kähler class on $X_{k}$ can be written as
$\omega_{t}]=\pi^{*}[\omega]-\sum_{i=1}^{k}t_{i}E_{i},$
where $E_{i}$ are the exceptional divisors. In these coordinates, the averaged scalar curvature in $\mathcal{J}_{X_{k}}$ is determined cohomologically, whereas the holomorphic $2$-systole $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])$ depends a priori on the effective curve classes on $X_{k}$. Thus, for a fixed blow-up type, the configuration dependence of $\mathcal{J}_{X_{k}}$ is encoded in the Mori cone. In particular, special configurations may produce additional effective curves, hence additional inequalities cutting out the admissible region in the Kähler cone. From the point of view of upper bounds for $\mathcal{J}_{X_{k}}$, this can only make the problem more restrictive. Accordingly, the largest possible supremum should be expected in the most generic configuration, when the blow-up points are as general as possible.
The difficulty is that the quantity $\operatorname{sys}^{\operatorname{hol}}_{2}([\omega_{t}])$ depends a priori on all effective curve classes on $X_{k}$, and when the Mori cone is not rational polyhedral this involves infinitely many possibilities. Our key observation is that one can nevertheless control $\mathcal{J}_{X_{k}}([\omega_{t}])$ using only finitely many numerical parameters extracted from the coefficients $t_{i}$, together with a mass-shift argument that identifies the worst possible distribution of these coefficients under fixed coarse data. This leads to an explicit finite-dimensional optimization problem over a coarse admissible region determined by finitely many effective curves. In this way, the geometric problem is reduced to a manageable numerical one. We carry this out first for blow-ups of $\mathbb{P}^{2}$, and then for blow-ups of ruled surfaces.
For non-rational ruled surfaces, we also give an independent analytic proof, inspired by Stern’s level-set method [22]. In the Kähler setting, the holomorphic fibration provides a natural replacement for the harmonic map used in the Riemannian argument. This yields the same sharp constant and the same rigidity statement as the algebro-geometric approach.
@@
It is useful to contrast the above discussion with the purely Riemannian PSC setting in higher dimensions. On a closed Riemannian manifold $(M^{n},g)$, the homological $k$-systole is only defined when $H_{k}(M;\mathbb{Z})\neq 0$, and even in that case a lower bound on scalar curvature does not in general force any upper bound for $\operatorname{sys}_{k}(g)$. A basic family of examples is
$M=S^{k}\times S^{n-k},\qquad g_{t}=t^{2}g_{S^{k}}\oplus g_{S^{n-k}},\qquad t\geq 1,$
where $g_{S^{m}}$ is the round metric of sectional curvature $1$. Then $\text{scal}(g_{t})$ stays uniformly positive, whereas the $k$-dimensional factor is stretched and $\operatorname{sys}_{k}(g_{t})\to\infty$ as $t\to\infty$. Thus positive scalar curvature alone does not control homological systoles in general.
By contrast, Kähler geometry provides additional structure. In basic Kähler examples, such as projective space, the relevant even-dimensional homology groups are already nontrivial, so the corresponding systolic questions are non-vacuous. More importantly for the present paper, on a compact Kähler surface the natural $2$-dimensional competitors are complex curves, and their areas are encoded cohomologically by the Kähler class.
In the PSC Kähler setting, the existence of such curves is moreover forced by the geometry itself: positive scalar curvature implies uniruledness, and hence the presence of many rational curves. This is the key difference from the general Riemannian situation. For PSC Kähler surfaces, one has canonical $2$-dimensional representatives and a rigid intersection-theoretic way to measure their areas. This is precisely what makes the holomorphic $2$-systole and the functional $\mathcal{J}_{X}$ the right objects for our problem.
@@
In Section 2 we review the basic notions used throughout the paper. We also show that, for projective Kähler surfaces with $c_{1}(X)\cdot[\omega]>0$, the functional $\mathcal{J}_{X}$ is well-defined and satisfies $\sup_{\mathcal{K}^{+}(X)}\mathcal{J}_{X}<\infty$.
@@
In Section 4 we study ruled surfaces from an algebro-geometric perspective. We first treat the non-rational ruled case, which is technically simpler in our approach, and show that $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 4\pi$, together with the corresponding rigid model. We then recall the geometry of Hirzebruch surfaces and their blow-ups, and adapt the mass-shift argument to the rational ruled case. This leads to the sharp bound $\min_{X}S(\omega)\cdot\operatorname{sys}_{2}(\omega)\leq 8\pi$ for all PSC Kähler surfaces fibred over $\mathbb{P}^{1}$, with rigidity characterized by the product Fubini-Study metric on $\mathbb{P}^{1}\times\mathbb{P}^{1}$.